# A Bayesian framework for active object recognition, pose estimation and shape transfer learning through touch

### v1

@@ III-A Problem Formulation
Let $\boldsymbol{D}$ denote the set of all observed tactile signals $\boldsymbol{d}$ and $\boldsymbol{X}$ be the set of all locations $\boldsymbol{x}$ where $\boldsymbol{D}$ are observed. Given tactile sensory data $\boldsymbol{D}$ at locations $\boldsymbol{X}$, the aim is to estimate the latent variables $\boldsymbol{z}$. Based on Bayes’ theorem, the posterior distribution on $\boldsymbol{z}$ can be calculated as follows,
$p(\boldsymbol{z}|\boldsymbol{D},\boldsymbol{X})=\frac{p(\boldsymbol{D}|\boldsymbol{z},\boldsymbol{X})p(\boldsymbol{z})}{\sum_{\boldsymbol{z}}p(\boldsymbol{D},\boldsymbol{z}|\boldsymbol{X})}\propto p(\boldsymbol{D}|\boldsymbol{z},\boldsymbol{X})p(\boldsymbol{z})$ (7)
where $p(\boldsymbol{z})$ is the prior joint distribution of object class and object pose, $p(\boldsymbol{D}|\boldsymbol{z},\boldsymbol{X})$ denotes the measurement likelihood function, which represents the likelihood of observing tactile sensory data $\boldsymbol{D}$ at locations $\boldsymbol{X}$ assuming the object class is $c$ and the pose is $\boldsymbol{p}$; $p(\boldsymbol{d}|\boldsymbol{z},\boldsymbol{x})$ is the observation likelihood function for a single data point. In this paper, $p(\boldsymbol{z})$ is chosen to be a uninformative uniform distribution $U$. For simplicity, in our formulation, each observation $\boldsymbol{d}$ at location $\boldsymbol{x}$ is assumed to be independent from other observations, given the latent variables $\boldsymbol{z}$ is known. This simplifies equation (7) into,
$p(\boldsymbol{D}|\boldsymbol{z},\boldsymbol{X})=\underset{i=1}{\prod^{O}}p(\boldsymbol{d}_{i}|\boldsymbol{z},\boldsymbol{x}_{i}),$ (8)
@@
As the proposed framework actively acquires data, at each time step $t$, new tactile data $\boldsymbol{D_{t}}$ are observed at locations $\boldsymbol{X}_{t}$. Let $T$ denote the total time steps passed, $\boldsymbol{D}_{1:T}$ represent all observed tactile data from time step $1$ to $T$ and $\boldsymbol{X}_{1:T}$ denote the locations where $\boldsymbol{D}_{1:T}$ are observed; then equation (7) can be rewritten as
$p(\boldsymbol{z}|\boldsymbol{D}_{1:T},\boldsymbol{X}_{1:T})=\frac{p(\boldsymbol{z},\boldsymbol{D}_{1:T}|\boldsymbol{X}_{1:T})}{\sum_{\boldsymbol{z}}p(\boldsymbol{z},\boldsymbol{D}_{1:T}|\boldsymbol{X}_{1:T})}.$ (9)
As we assume independence between different observations given $\boldsymbol{z}$, the joint distribution can be calculated as
$p(\boldsymbol{z},\boldsymbol{D}_{1:T}|\boldsymbol{X}_{1:T})$ $=p(\boldsymbol{z})p(\boldsymbol{D}_{1:T-1}|\boldsymbol{z},\boldsymbol{X}_{1:T-1})p(\boldsymbol{D}_{T}|\boldsymbol{z},\boldsymbol{X}_{T})$ (10)
$=p(\boldsymbol{z},\boldsymbol{D}_{1:T-1}|\boldsymbol{X}_{1:T-1})p(\boldsymbol{D}_{T}|\boldsymbol{z},\boldsymbol{X}_{T}).$
One can obtain the MLE $\boldsymbol{z^{*}}$ of the object pose and object class given observed data as follows,
$\boldsymbol{z^{*}}:=\operatorname*{argmax}_{\boldsymbol{z}}(p(\boldsymbol{z}|\boldsymbol{D}_{1:T},\boldsymbol{X}_{1:T}))\$ (11)
To decide if the object is known or unknown, one can track the consistency between the $\boldsymbol{z}^{*}$ and the observed data $\boldsymbol{D}_{1:T}$ by calculating $p(\boldsymbol{D}_{1:T}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:T})$ using equation (8). In the following sections, $p(\boldsymbol{D}_{1:T}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:T})$ is referred to as MLE model evidence. A low value for $p(\boldsymbol{D}_{1:T}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:T})$, when the object surface is fully explored, would imply the test object is unknown.
III-C Measurement Likelihood Function
The measurement likelihood function is defined based on the difference between the predictive signed distance values, normal vectors and the observed values. The observed signed distance value $d_{s}$ at a contact location $\boldsymbol{x}$ is always regarded as zero.
If we assume that each tactile observation $\boldsymbol{d}$ at location $\boldsymbol{x}$ follows a Gaussian distribution,
$\boldsymbol{d}\sim N(\boldsymbol{\widehat{d}},\boldsymbol{\Sigma})$ (12)
$\boldsymbol{\Sigma}=\text{diag}\left(\left[\frac{1}{2\sigma_{d}^{2}},\frac{1}{2\sigma_{n}^{2}},\frac{1}{2\sigma_{n}^{2}},\frac{1}{2\sigma_{n}^{2}}\right]\right).$ (13)
In other words, the measurement likelihood function satisfies:
$p(\boldsymbol{d}|\boldsymbol{z},\boldsymbol{x})=\frac{1}{Z}\exp\left(\boldsymbol{\widehat{d}}-\boldsymbol{d}\right)^{T}\boldsymbol{\Sigma}\left(\boldsymbol{\widehat{d}}-\boldsymbol{d}\right)$ (14)
$Z=(2\pi)^{2}\sqrt{\det(\boldsymbol{\Sigma})}$ (15)
$-\ln p(\boldsymbol{d}|\boldsymbol{z},\boldsymbol{x})\propto\left(\boldsymbol{\widehat{d}}-\boldsymbol{d}\right)^{T}\boldsymbol{\Sigma}\left(\boldsymbol{\widehat{d}}-\boldsymbol{d}\right),$ (16)
where diag is an operator to form a diagonal matrix from a vector, Z is the normalization constant, $\sigma_{d}$ and $\sigma_{n}$ denote the standard deviations for the signed distance function observation and surface normal observation respectively.
III-D Negative Information Update
Equation (14) assumes the observed data takes on precise values. However, when the tactile sensor is not in contact with the object, it implies that, at the sensor location $\boldsymbol{x}$, $d_{s}\in(0,+\infty)$, and though the normal vector can still be predicted at $\boldsymbol{x}$, it cannot be observed from tactile sensors as no contact force would be sensed in a real setup.
@@ III-D Negative Information Update
$p(d_{s}>0|\boldsymbol{z},\boldsymbol{x})=\frac{1}{2}\left(1-\mathbf{erf}\left(\frac{-f(\boldsymbol{x},\boldsymbol{z})}{(\sqrt{2}\sigma_{d}}\right)\right).$ (18)
where $\mathbf{erf}$ denotes the Gauss error function. By substituting equation (17) and equation (18) into equation (8) the framework can update it’s belief when no contact is observed at a location $\boldsymbol{x}$.
III-E Particle Filter (PF) for Object Recognition and Pose Estimation
A particle filter (PF) [5] is adopted to estimate the joint distribution of the object’s class and pose online based on equations (11)-(16). The core idea behind particle filters is to use weighted discrete samples (particles) to approximate the posterior distribution of interest based on importance sampling, i.e.,
$p(\boldsymbol{z}|\boldsymbol{D}_{1:t},\boldsymbol{X}_{1:t})=\sum_{j=1}^{N}w_{j,t}\;\delta(\boldsymbol{z}-\boldsymbol{z}_{j,t}),$ (19)
where the subscript $j,t$ denotes the index of a particle in the PF at time step $t$, and $\delta$ denotes the Dirac delta function, $N$ denotes total number of particles. The unnormalized weight of a particle $w_{j,t}$ can be calculated as
$w_{j,t}$ $\propto p(\boldsymbol{z}_{j,t}|\boldsymbol{D}_{1:t},\boldsymbol{X}_{1:t})$ (20)
$\propto p(\boldsymbol{z}_{j,t},D_{t}|\boldsymbol{D}_{1:t-1},\boldsymbol{X}_{1:t})$
$=p(\boldsymbol{z}_{j,t}|\boldsymbol{D}_{1:t-1},\boldsymbol{X}_{1:t-1})\;p(\boldsymbol{D}_{t}|\boldsymbol{z}_{j,t},\boldsymbol{X}_{t})$
$=\sum_{\boldsymbol{z}_{j,t-1}}(\overline{w}_{j,t-1}\;p(\boldsymbol{z}_{j,t}|\boldsymbol{z}_{j,t-1},\boldsymbol{D}_{1:t-1},\boldsymbol{X}_{1:t-1}))\times$
$\;\;\;\;\;\;\;p(\boldsymbol{D}_{t}|\boldsymbol{z}_{j,t},\boldsymbol{X}_{t}).$
@@ III-E Particle Filter (PF) for Object Recognition and Pose Estimation
$p(\boldsymbol{z}_{j,t}|\boldsymbol{z}_{j,t-1},\boldsymbol{D}_{1:t-1},\boldsymbol{X}_{1:t-1})$ $=\delta(\boldsymbol{z}_{j,t}-\boldsymbol{z}_{j,t-1}),$ (21)
$\boldsymbol{z}_{j,t}=\boldsymbol{z}_{j,t-1}.$ (22)
Noteworthy, $p(\boldsymbol{z}_{j,t}|\boldsymbol{z}_{j,t-1},\boldsymbol{D}_{1:t-1},\boldsymbol{X}_{1:t-1})$ can be replaced by a motion estimator if a motion model is available. With equation (21) and equation (22), it follows
$w_{j,t}$ $=\sum_{\boldsymbol{z}_{j,t-1}}(\overline{w}_{j,t-1}\;\delta(\boldsymbol{z}_{j,t}-\boldsymbol{z}_{j,t-1}))\;p(\boldsymbol{D}_{t}|\boldsymbol{z}_{j,t},\boldsymbol{X}_{t})$ (23)
$=\overline{w}_{j,t-1}\;p(\boldsymbol{D}_{t}|\boldsymbol{z}_{j,t},\boldsymbol{X}_{t})$
$\overline{w}_{j,t}=\frac{w_{j,t}}{\sum_{j=1}^{N}w_{j,t}}\;$ (24)
where $\overline{w}_{j,t}$ is the normalized weight of particle $\boldsymbol{z}_{j,t}$. Note that the recursive formula of the particle weight update is consistent with equation(10).
As the particle filter needs to estimate the joint distribution of the object’s class and its 6D pose, the number of particles required to cover the space sufficiently can be calculated as $n\times m^{6}$, where $n$ denotes the number of known object classes, and $m$ denotes the resolution of the discretization at each continuous dimension.
To enhance the sample efficiency of the particle filter and have better coverage on the high-density region of the posterior distribution, we propose to sample new particles based on newly observed data at each time step and mix them with existing particles. The rotation and translation invariant point-pair feature proposed in [4] is adopted for this purpose. For every two data points, the point-pair feature can be calculated. For more details about the alignment process, please refer to [4].
Figure 2: An example of sampling particles with point-pair features. For each observed data point pair, $\theta_{i,j},\theta_{i,d},\theta_{j,d}$ and $||\boldsymbol{x}_{i,j}||_{2}$ are calculated as the point pair features. $\boldsymbol{x}_{i,j}$ denotes the vector $\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$. $\theta_{i,j}$ ,$\theta_{i,d}$, $\theta_{j,d}$ are the angle between normal vectors of $\boldsymbol{n}_{i}$ and $\boldsymbol{n}_{j}$, the angle between $\boldsymbol{n}_{i}$, and $\boldsymbol{x}_{i,j}$, the angles between $\boldsymbol{n}_{j}$ and $\boldsymbol{x}_{i,j}$ respectively. Point pairs on all known models with similar features are extracted. Finally, by aligning known model point pairs with the observed point pair, possible combinations of the object class and object pose that match the observed point pair are found. $c_{s}$, $c_{e}$, $c_{r}$ represent three object classes, namely square, ellipse, and rectangular respectively. $\boldsymbol{p_{s}^{0}}$, $\boldsymbol{p_{e}^{0}}$, $\boldsymbol{p_{r}^{0}}$ are the original object pose (by convention set to $[0,0,0,0,0,0]$), whereas $\boldsymbol{p_{s}^{{}^{\prime}}}$, $\boldsymbol{p_{e}^{{}^{\prime}}}$, $\boldsymbol{p_{r}^{{}^{\prime}}}$ are the pose after the alignment.
@@
For each known object, 200 feature points are sampled using Poisson Disc Sampling[31], resulting in $200*199=39800$ point pairs. The point-pair features for point pairs on known models are pre-computed, discretized and stored in a hash table for quick lookup at runtime. Each key in the table corresponds to a discretized feature value, and under the key is a list of point pairs on the known models that yield this discretized feature value. Each element in the list contains the indices of two points and their corresponding object class. Upon a newly obtained contact point, new point-pair features between the new contact point and previous contact points are calculated, discretized, and used to look for point pairs on the known models with the same point-pair features using the hash table. The discretization is done by rounding $||\boldsymbol{x}_{i,j}||_{2}$ to the closest one digit decimal and $\theta_{i,j}$ ,$\theta_{i,d}$, $\theta_{j,d}$ to the closest multiple of $12$ degrees. Importantly, the resolution of the discretization should be determined by the estimated noise level of the tactile signal, such that matches can be found during runtime. In the simulation, the noise level of the signal is known therefore it can be set directly. The lookup table process is used to find possible correspondence point pairs on known object models.
@@
The lookup table and alignment processes together can be viewed as sampling particles from the mixture of posterior distributions $p(\boldsymbol{z}|\boldsymbol{x}_{i},\boldsymbol{n}_{i},\boldsymbol{x}_{j},\boldsymbol{n}_{j})$ conditioned on each observed point pair. The core idea is that the sampling procedure can concentrate new particles in high probability density regions of the true posterior distribution which the PF aims to approximate.
@@
As tactile data points are obtained sequentially in this study, at each time step $t$, new particles are sampled based on the new point-pair features. At time step $t$, if pairing the new contact point with all previously observed contact points $t-1$, the number of new point pairs is $t-1$, taking into the order of the points, in total $2(t-1)$ pairs need to be considered, which is undesirable as $t$ increases. Additionally, to assign proper weights to new particles, they need to be evaluated on previous observations and it is computationally expensive to evaluate all new particles on all previous observations.
@@
To this end, we propose a method to select a fixed number of contact locations $\boldsymbol{X}_{s}$ from $\boldsymbol{X}_{1:t-1}$ at time step $t$, to form a fixed number of new point-pair features, while evaluating the new particles sampled from these features on the observed data $\boldsymbol{D}_{s}$ at $\boldsymbol{X}_{s}$, as follows. A subset of previous observed tactile data $\boldsymbol{D}_{1:t-1}$ at $\boldsymbol{X}_{1:t-1}$ is selected to determine the relative weights among new particles within each object class. All previous contact locations in $\boldsymbol{X}_{1:t-1}$ are first sorted based on their distance to the new observed contact location, and then the sorted contact locations are divided into $n_{s}$ adjacent segments with the same size. The first point within each segment constitutes $\boldsymbol{X}_{s}$. The tactile data observed at $\boldsymbol{X}_{s}$ is selected as $\boldsymbol{D}_{s}$. As a result, at each time step, a fixed number, $2\times n_{s}$ of new point-pairs are considered, as the order of points matters due to discretization. A fixed number $n_{s}$ of points are used to calculate the relative weights. This choice is incentivized by locations with richer distance profiles that will capture the global context of the object better compared to using a temporal sliding window as proposed in [10].
@@
The derivation of the weight approximation scheme goes as follows. Let us denote the weight of new particles $\boldsymbol{z}^{{}^{\prime}}_{c,k,t}$ with class $c$ by $w^{{}^{\prime}}_{c,k,t}$ and the highest weight among them by $w^{{}^{\prime}}_{c,*,t}$. $w^{{}^{\prime}}_{c,*,t}$ need to satisfy the following equation:
$\frac{w^{{}^{\prime}}_{c,k,t}}{w^{{}^{\prime}}_{c,*,t}}=\frac{p(\boldsymbol{D}_{s}|\boldsymbol{z}^{{}^{\prime}}_{c,k,t},\boldsymbol{X}_{s})}{p(\boldsymbol{D}_{s}|\boldsymbol{z}^{{}^{\prime}}_{c,*,t},\boldsymbol{X}_{s})}.$ (25)

### v3

@@ III-A4 Bayesian Inference of Object Class and Object Pose
Let $\mathbf{D}$ denote the set of all observed tactile signals $\mathbf{d}$ and $\mathbf{X}$ be the set of all locations $\mathbf{x}$ where $\mathbf{D}$ are observed. Given tactile sensory data $\mathbf{D}$ at locations $\mathbf{X}$, the goal is to estimate the latent variables $\mathbf{z}$. Based on Bayes’ theorem, the posterior distribution on $\mathbf{z}$ can be calculated as follows,
$p(\mathbf{z}|\mathbf{D},\mathbf{X})=\frac{p(\mathbf{D}|\mathbf{z},\mathbf{X})p(\mathbf{z})}{\sum_{\mathbf{z}}p(\mathbf{D},\mathbf{z}|\mathbf{X})}\propto p(\mathbf{D}|\mathbf{z},\mathbf{X})p(\mathbf{z})$ (8)
where $p(\mathbf{z})$ is the prior joint distribution of object class and object pose when no data is present, which is chosen to be a uninformative uniform distribution $U$ in this study; $p(\mathbf{D}|\mathbf{z},\mathbf{X})$ denotes the likelihood of observing tactile sensory data $\mathbf{D}$ at locations $\mathbf{X}$ assuming the object class is $c$ and the pose is $\mathbf{p}$. The summation operation $\sum_{\mathbf{z}}$ is overloaded to represent the summation over $c$ and integrals over $\mathbf{p}$.
As the proposed framework actively acquires data, at each time step $t$, new tactile data $\mathbf{D_{t}}$ are observed at locations $\mathbf{X}_{t}$. Let $T$ denote the total time steps passed, $\mathbf{D}_{1:T}$ represent all observed tactile data from time step $1$ to $T$, and $\mathbf{X}_{1:T}$ denote the locations where $\mathbf{D}_{1:T}$ are observed; then (8) can be rewritten as
$p(\mathbf{z}|\mathbf{D}_{1:T},\mathbf{X}_{1:T})=\frac{p(\mathbf{z},\mathbf{D}_{1:T}|\mathbf{X}_{1:T})}{\sum_{\mathbf{z}}p(\mathbf{z},\mathbf{D}_{1:T}|\mathbf{X}_{1:T})}.$ (9)
By assuming independence between different observations given $\mathbf{z}$, the joint distribution can be calculated as
$p(\mathbf{z},\mathbf{D}_{1:T}|\mathbf{X}_{1:T})$ $=p(\mathbf{z})p(\mathbf{D}_{1:T-1}|\mathbf{z},\mathbf{X}_{1:T-1})p(\mathbf{D}_{T}|\mathbf{z},\mathbf{X}_{T})$ (10)
$=p(\mathbf{z},\mathbf{D}_{1:T-1}|\mathbf{X}_{1:T-1})p(\mathbf{D}_{T}|\mathbf{z},\mathbf{X}_{T}).$
Following the aforementioned independence assumption, $p(\mathbf{D}_{t}|\mathbf{z},\mathbf{X}_{t})$ can be further decomposed into
$p(\mathbf{D}_{t}|\mathbf{z},\mathbf{X}_{t})=\underset{i=1}{\prod^{O}}p(\mathbf{d}_{i,t}|\mathbf{z},\mathbf{x}_{i,t}).$ (11)
@@
III-A5 Measurement Likelihood Function
Assuming each tactile observation $\mathbf{d}$ at a location $\mathbf{x}$ follows a Gaussian distribution $\mathbf{d}\sim N(\mathbf{\widehat{d}},\boldsymbol{\Sigma})$, then the measurement likelihood function can be written as
$p(\mathbf{d}|\mathbf{z},\mathbf{x})=\frac{1}{Z}\exp\left[\left(\mathbf{\widehat{d}}-\mathbf{d}\right)^{T}\boldsymbol{\Sigma}\left(\mathbf{\widehat{d}}-\mathbf{d}\right)\right],$ (12)
$Z=(2\pi)^{2}\sqrt{\det(\boldsymbol{\sigma})},$ (13)
$\boldsymbol{\sigma}=\text{diag}\left(\left[\frac{1}{2\sigma_{d}^{2}},\frac{1}{2\sigma_{n}^{2}},\frac{1}{2\sigma_{n}^{2}},\frac{1}{2\sigma_{n}^{2}}\right]\right),$ (14)
where $Z$ is the normalization constant, diag is an operator to form a diagonal matrix from a vector, $\sigma_{d}$ and $\sigma_{n}$ denote the standard deviations for the signed distance observation and surface normal vector observation, respectively.
III-A6 Negative Information Update
Eqn. (12) assumes the observed data $\mathbf{d}$ takes on precise values. However, when the tactile sensor is not in contact with the object, it indicates that $d_{s}\in(0,+\infty)$ at location $\mathbf{x}$, and the surface normal vector will not be taken into account for the likelihood calculation. This entails,
@@ III-A6 Negative Information Update
$p(d_{s}>0|\mathbf{z},\mathbf{x})=\frac{1}{2}\left(1-\mathbf{erf}\left(\frac{-f(\mathbf{x},\mathbf{z})}{\sqrt{2}\sigma_{d}}\right)\right).$ (16)
where $\mathbf{erf}$ denotes the Gauss error function. By substituting (15)(16) into (11), the framework can update its belief using non-contact points.
III-A7 MAP Model Evidence
One can obtain the MAP combination of the object pose and object class, denoted by $\mathbf{z^{*}}$, as follows,
$\mathbf{z^{*}}:=\operatorname*{\text{arg max}}_{\mathbf{z}}(p(\mathbf{z}|\mathbf{D}_{1:T},\mathbf{X}_{1:T}))\$ (17)
The framework tracks the fitness between the $\mathbf{z}^{*}$ and the observed data $\mathbf{D}_{1:T}$ by calculating $p(\mathbf{D}_{1:T}|\mathbf{z}^{*},\mathbf{X}_{1:T})$ using (11). In the following sections, $p(\mathbf{D}_{1:T}|\mathbf{z}^{*},\mathbf{X}_{1:T})$ is referred to as MAP model evidence. A low value for $p(\mathbf{D}_{1:T}|\mathbf{z}^{*},\mathbf{X}_{1:T})$ would imply the test object is novel. The threshold for classifying the test object as a novel object will be derived in another subsection later.
III-B Customized Particle Filter for Object Recognition and Pose Estimation
Combining (9)-(16), one can update the posterior distribution $p(\mathbf{z}|\mathbf{D}_{1:t},\mathbf{X}_{1:t})$ recursively for each time step $t$ in theory. However, the term $\sum_{\mathbf{z}}p(\mathbf{z},\mathbf{D}_{1:T}|\mathbf{X}_{1:T})$ in (9) is intractable in practice. To address this issue, a particle filter (PF) is adopted to estimate the joint distribution of the object class and object pose continuously. The core idea behind particle filters is to use weighted discrete samples (particles) to approximate the posterior distribution of interest based on importance sampling, i.e.,
$p(\mathbf{z}|\mathbf{D}_{1:t},\mathbf{X}_{1:t})=\sum_{j=1}^{N}\overline{w}_{j,t}\;\delta(\mathbf{z}-\mathbf{z}_{j,t}),$ (18)
where the subscript $j,t$ denotes the index of a particle in the PF at time step $t$, $\overline{w}_{j,t}$ denotes the normalized weight of the particle $\mathbf{z}_{j,t}$, $\delta$ denotes the Dirac delta function, and $N$ denotes total number of particles. The unnormalized weight $w_{j,t}$ of the particle $\mathbf{z}_{j,t}$ can be calculated as
$w_{j,t}$ $\propto p(\mathbf{z}_{j,t}|\mathbf{D}_{1:t},\mathbf{X}_{1:t})$ (19)
$\propto p(\mathbf{z}_{j,t},\mathbf{D}_{t}|\mathbf{D}_{1:t-1},\mathbf{X}_{1:t})$
$=p(\mathbf{z}_{j,t}|\mathbf{D}_{1:t-1},\mathbf{X}_{1:t-1})\;p(\mathbf{D}_{t}|\mathbf{z}_{j,t},\mathbf{X}_{t})$
$=\sum_{\mathbf{z}_{j,t-1}}(\overline{w}_{j,t-1}\;p(\mathbf{z}_{j,t}|\mathbf{z}_{j,t-1},\mathbf{D}_{1:t-1},\mathbf{X}_{1:t-1}))\times$
$\;\;\;\;\;\;\;p(\mathbf{D}_{t}|\mathbf{z}_{j,t},\mathbf{X}_{t}).$
@@ III-B Customized Particle Filter for Object Recognition and Pose Estimation
$p(\mathbf{z}_{j,t}|\mathbf{z}_{j,t-1},\mathbf{D}_{1:t-1},\mathbf{X}_{1:t-1})$ $=\delta(\mathbf{z}_{j,t}-\mathbf{z}_{j,t-1}),$ (20)
$\mathbf{z}_{j,t}=\mathbf{z}_{j,t-1}.$ (21)
Noteworthy, $p(\mathbf{z}_{j,t}|\mathbf{z}_{j,t-1},\mathbf{D}_{1:t-1},\mathbf{X}_{1:t-1})$ can be replaced by a motion estimator if a motion model is available. With (20) and (21), it follows
$w_{j,t}$ $=\sum_{\mathbf{z}_{j,t-1}}(\overline{w}_{j,t-1}\;\delta(\mathbf{z}_{j,t}-\mathbf{z}_{j,t-1}))\;p(\mathbf{D}_{t}|\mathbf{z}_{j,t},\mathbf{X}_{t})$ (22)
$=\overline{w}_{j,t-1}\;p(\mathbf{D}_{t}|\mathbf{z}_{j,t},\mathbf{X}_{t}),$
$\overline{w}_{j,t}=\frac{w_{j,t}}{\sum_{j=1}^{N}w_{j,t}}.$ (23)
where $\overline{w}_{j,t}$ is the normalized weight of particle $\mathbf{z}_{j,t}$. Using (23) and (23), one can update each particle’s weight recursively at each time step.
As the particle filter needs to estimate the joint distribution of the object’s class and its 6-DOF pose, the number of particles required to cover the space sufficiently can be calculated as $n\times(m^{6})$, where $n$ denotes the number of known object classes, and $m$ denotes the resolution of the discretization at each continuous dimension.
To enhance the sample efficiency of the particle filter and have better coverage on the high-density region of the posterior distribution, we propose to sample new particles based on newly observed data at each time step and mix them with existing particles. The rotation and translation invariant point-pair feature proposed in [9] is adopted for this purpose. For every two data points, the point-pair feature can be calculated.
Figure 2: An example of sampling particles with point-pair features. For each observed data point pair, $\theta_{i,j},\theta_{i,d},\theta_{j,d}$ and $||\mathbf{x}_{i,j}||_{2}$ are calculated as the point pair features. $\mathbf{x}_{i,j}$ denotes the vector $\mathbf{x}_{j}-\mathbf{x}_{i}$. $\theta_{i,j}$ ,$\theta_{i,d}$, $\theta_{j,d}$ are the angle between normal vectors of $\mathbf{n}_{i}$ and $\mathbf{n}_{j}$, the angle between $\mathbf{n}_{i}$, and $\mathbf{x}_{i,j}$, the angles between $\mathbf{n}_{j}$ and $\mathbf{x}_{i,j}$ respectively. Point pairs on all known models with similar features are extracted. Finally, by aligning known model point pairs with the observed point pair, possible combinations of the object class and object pose that match the observed point pair are found. $c_{s}$, $c_{e}$, $c_{r}$ represent three object classes, namely square, ellipse, and rectangular respectively. $\mathbf{p_{s}^{0}}$, $\mathbf{p_{e}^{0}}$, $\mathbf{p_{r}^{0}}$ are the original object pose, by convention set to $[0,0,0,0,0,0]$, whereas $\mathbf{p_{s}^{{}^{\prime}}}$, $\mathbf{p_{e}^{{}^{\prime}}}$, $\mathbf{p_{r}^{{}^{\prime}}}$ are the pose after the alignment.
@@
For each known object, 200 feature points are sampled using Poisson Disc Sampling [29], resulting in $200*199=39800$ point pairs. The point-pair features for point pairs on known models are pre-computed, discretized and stored in a hash table for quick look-up at runtime. Each key in the table corresponds to a discretized feature value, and under the key is a list of point pairs on the known models that yield this discretized feature value. Each element in the list contains the indices of two points and their corresponding object class. Upon a newly obtained contact point, new point-pair features between the new contact point and previous contact points are calculated, discretized, and used to look for point pairs on the known models with the same point-pair features using the hash table. The discretization is done by rounding $||\mathbf{x}_{i,j}||_{2}$ to the closest one digit decimal and $\theta_{i,j}$ ,$\theta_{i,d}$, $\theta_{j,d}$ to the closest multiple of $12$ degrees. Importantly, the resolution of the discretization should be determined by the estimated noise level of the tactile signal, such that matches can be found during runtime. In the simulation, the noise level of the signal is known, therefore it can be set directly. The look-up table process is used to find possible correspondence point pairs on known object models.
@@
The look-up table and alignment processes together can be viewed as sampling particles from the mixture of posterior distributions $p(\mathbf{z}|\mathbf{x}_{i},\mathbf{n}_{i},\mathbf{x}_{j},\mathbf{n}_{j})$ conditioned on each observed point pair. The core idea is that the sampling procedure can concentrate new particles in high probability density regions of the true posterior distribution which the PF aims to approximate.
@@
As tactile data points are obtained sequentially in this study, at each time step $t$, new particles are sampled based on the new point-pair features. At time step $t$, if pairing the new contact point with all previously observed contact points $t-1$, the number of new point pairs is $t-1$. Taking into account the order of the points, in total $2(t-1)$ pairs need to be considered, which is undesirable as $t$ increases.
@@
Additionally, to assign proper weights to new particles, they need to be evaluated on previous observations and it is computationally expensive to evaluate all new particles on all previous observations. To this end, we propose a method to select a fixed number of contact locations $\mathbf{X}_{s}$ from $\mathbf{X}_{1:t-1}$ at time step $t$, to form a fixed number of new point-pair features, while evaluating the new particles sampled from these features on the observed data $\mathbf{D}_{s}$ at $\mathbf{X}_{s}$, as follows. A subset of previous observed tactile data $\mathbf{D}_{1:t-1}$ at $\mathbf{X}_{1:t-1}$ is selected to determine the relative weights among new particles within each object class. All previous contact locations in $\mathbf{X}_{1:t-1}$ are first sorted based on their distance to the new observed contact location, and then the sorted contact locations are divided into $n_{s}$ adjacent segments with the same size. The first point within each segment constitutes $\mathbf{X}_{s}$. The tactile data observed at $\mathbf{X}_{s}$ is selected as $\mathbf{D}_{s}$. As a result, at each time step, a fixed number, $2\times n_{s}$ of new point-pairs are considered, as the order of points matters due to discretization. A fixed number $n_{s}$ of points are used to calculate the relative weights. This choice is incentivized by locations with richer distance profiles that will capture the global context of the object better compared to using a temporal sliding window as proposed in [6].
@@
The derivation of the weight approximation scheme goes as follows. Let us denote the weight of new particles $\mathbf{z}^{{}^{\prime}}_{c,k,t}$ with class $c$ by $w^{{}^{\prime}}_{c,k,t}$ and the highest weight among them by $w^{{}^{\prime}}_{c,*,t}$. $w^{{}^{\prime}}_{c,*,t}$ need to satisfy the following equation:
$\frac{w^{{}^{\prime}}_{c,k,t}}{w^{{}^{\prime}}_{c,*,t}}=\frac{p(\mathbf{D}_{s}|\mathbf{z}^{{}^{\prime}}_{c,k,t},\mathbf{X}_{s})}{p(\mathbf{D}_{s}|\mathbf{z}^{{}^{\prime}}_{c,*,t},\mathbf{X}_{s})}.$ (24)